# Occupancy-Predictive Path Planning for Autonomous Surface Vehicle Navigation in Ice-Covered Water

### v1

@@ I Introduction
While path planning for ASV ice navigation has been extensively studied, much of the existing literature focuses on computing a path on a global scope [8, 9, 10]. For instance, the planned paths in [8] and [9] span hundreds of kilometers while the paths considered in [10] take days to traverse. While these works demonstrated good performance, local planners are required to achieve full ship autonomy. In ASV local planning, however, most work aims to compute a collision-free path [4, 5, 6, 11].
@@ III-C2 Loss Functions
To guide the learning process toward more physically-plausible predictions, we present a novel physics-based loss tailored to our occupancy formulation. The core assumption is that the total mass of the ice floes is conserved before and after the collisions from the ship executing a maneuver. Let $m_{p}$ represent the mass of an ice floe $p\in W_{obs}$, and let $W_{obs}^{i}=\{p^{i}_{1},\ldots p^{i}_{m}\}$ and $W_{obs}^{f}=\{p^{f}_{1},\ldots p^{f}_{n}\}$ be the set of obstacle ice floes before and after ship executing a motion primitive, respectively. By conservation of mass, we have
@@
Similar to [3], we assume that the ice floes are 3D polygons with uniform density $\rho$ and thickness $d$. Since the mass of an ice floe is the product of its density, thickness, and surface area, the total surface area of the ice flows is conserved. However, each grid in the occupancy map is a ratio between the obstacle-occupied area and the grid area. Thus, the conservation of surface area implies the conservation of the sum of occupancy values in the map (see the extended version [23] for a complete derivation):
$\textup{Sum}(O)$ $=\textup{Sum}(O^{\prime}).$ (5)
@@ IV-B Collision Cost with Occupancy
We begin by presenting our empirical analysis on how the occupancy change, denoted as $\textup{diff}(O,O^{\prime})$, that results from a ship motion primitive relates to the ship’s kinetic energy loss, collision impulse and approximated work done ($W_{\text{approx}}$ described in Sec. V-A) during the execution of the motion primitive. For this analysis, we use occupancy maps that match the input/output dimensions of our learned predictor.
@@
$c_{\pi}=\textup{diff}(O,O^{\prime}).$ (7)
@@
$C(\Pi)=\sum_{\pi\in\Pi}\text{diff}(O_{\pi},O^{\prime}_{\pi}).$ (8)
@@ IV-C A* Search with Occupancy Estimation
During each iteration, the current node $n$ is expanded to its neighbor $n^{\prime}$ via a feasible motion primitive $\pi$, denoted by the overloaded $+$ operator at Line 11. A local occupancy $O$ is cropped from the global occupancy $O^{\textup{global}}$ based on the ship pose at node $n$ (Line 13) and is then used for occupancy prediction (Line 14). The edge cost of $\pi$ is then computed as a linear combination of the occupancy-difference collision cost discussed in IV-B scaled by $\alpha$ and the motion primitive length $d(\pi)$ (Line 15). The predicted local occupancy $O^{\prime}$ is stitched back for an updated global occupancy map $O^{\prime\textup{global}}$ (Line 16), which is then cached if the neighbor $n^{\prime}$ is visited for the first time or revisited with a better path (Lines 22). In a future iteration when the node $n^{\prime}$ is being expanded, the global occupancy at $n^{\prime}$ predicted previously is retrieved for further predictions (Line 9).
@@ 18                   Update $n^{\prime}$ parent pointer
Runtime: While Alg. 1 has the same structure as $A^{*}$ or Dijkstra’s algorithm, the bottleneck operation is the neural network prediction in Line 14, rather than the priority queue operation. Note that in Alg. 1, each node is expanded only once. Let $|E|$ be the number of edges in the expanded search graph from Alg. 1, it follows that the runtime is $O(|E|P)$, where $P$ is the time for each forward pass of the predictor. A more thorough discussion of runtime is provided in [23].
Performance: Note that by considering obstacle motions, the path finding problem lacks optimal substructure. Further, the general problem of navigation among movable obstacle is known to be NP-hard [25]. Consequently, Alg. 1 lacks a theoretical optimality guarantee. Hence we present the following constant factor guarantee on Alg. 1.
@@
Let $I$ be a problem instance, $\textup{OPT}(I)$ be the optimal path and $\textup{Alg}(I)$ be the path returned by Alg. 1 assuming perfect predictions. Let $l_{\min}$ be the distance of the shortest action primitive, the cost of $\textup{Alg}(I)$ is bounded as
$u(\textup{Alg}(I))\leq(1+\frac{\alpha}{l_{\min}})\cdot u(\textup{OPT}(I)).$ (9)
The proof of Thm. 1 is found in the extended version [23].
V Evaluation
@@
V-A Simulation Setup
We conduct evaluations on the autonomous ship-ice navigation simulator from [3]. The setup is consistent with the experimental platform at the National Research Council Canada ice tank in St. John’s, NL, featuring a 1:45 scale model vessel. Tests are conducted at 20%, 30%, 40%, and 50% concentrations, with 200 randomly generated environments per concentration. Fig. 1 (left) shows a 40% concentration ice field. We evaluate all four planners on the same set of trials, giving a total of $4\,\times 200\,\times 4\,=3200\,\text{experiments}$.
@@ V-B Simulation Results
V-C Physical Testbed Setup
@@
V-D Physical Testbed Evaluation
We perform testbed evaluations in 20%, 30%, and 40% concentration environments. For each concentration, 40 trials are performed for each planner, giving a total of $3\,\times 40\,\times 4\,=480\,\text{experiments}$. $W_{\text{approx}}$ and travel distance are used as metrics to quantify collision minimization and travel efficiency, respectively. Demonstrations of the testbed setup and experiments can be found in the supplemental video.

### v2

# Autonomous Navigation in Ice-Covered Waters with Learned Predictions on Ship-Ice Interactions

@@ I Introduction
While path planning for ASV ice navigation has been extensively studied, much of the existing literature focuses on computing a path on a global scope [8, 9, 10]. For instance, the planned paths in [8] and [9] span hundreds of kilometers while the paths considered in [10] take days to traverse. While these works demonstrated good performance, local planners are required to achieve full ship autonomy. In ASV local planning, however, most work aims to compute a collision-free path [4, 5, 6, 11], which does not generalize to environments where a collision-free path does not exist.
@@ III-C2 Loss Functions
To guide the learning process toward more physically-plausible predictions, we present a novel physics-based loss tailored to our occupancy formulation. The core assumption is that the total mass of the ice floes is conserved before and after the collisions from the ship executing a maneuver. Let $m_{p}$ represent the mass of an ice floe $p$, and let $W_{obs}^{i}=\{p^{i}_{1},\ldots p^{i}_{m}\}$ and $W_{obs}^{f}=\{p^{f}_{1},\ldots p^{f}_{n}\}$ be the sets of obstacle ice floes before and after ship executing a motion primitive, respectively. By conservation of mass, we have
@@
Similar to [3], we assume that the ice floes are 3D polygons with uniform density $\rho$ and thickness $d$. It follows that the mass of an ice floe is $m_{p}=\sigma_{p}\rho d$, where $\sigma_{p}$ denotes the surface area of an ice floe $p$. We rewrite Eq. 4 as
$\begin{split}\sum_{p^{i}\in W_{obs}^{i}}\sigma_{p^{i}}\rho d=\sum_{p^{f}\in W_{obs}^{f}}\sigma_{p^{f}}\rho d\\ \sum_{p^{i}\in W_{obs}^{i}}\sigma_{p^{i}}=\sum_{p^{f}\in W_{obs}^{f}}\sigma_{p^{f}}.\end{split}$ (5)
Recall that each grid in our occupancy map is a ratio between the obstacle-occupied area and the grid area. Let $s_{g}$ denote the grid size of the occupancy map, and let $o_{p}$ be the total occupancy value the ice floe $p$ contributes to the occupancy map, it follows that $\sigma_{p}=s_{g}o_{p}$. Further, note that the sum of all ice floes’ occupancy $o_{p}$ is simply the sum of occupancy values in the occupancy map, namely $\sum_{p^{i}\in W_{obs}^{i}}o_{p^{i}}=\textup{Sum}(O)$ and $\sum_{p^{f}\in W_{obs}^{f}}o_{p^{f}}=\textup{Sum}(O^{\prime})$. With these observations, Eq. 4 can be rewritten as
$\begin{split}\sum_{p^{i}\in W_{obs}^{i}}o_{p^{i}}s_{g}=\sum_{p^{f}\in W_{obs}^{f}}o_{p^{f}}s_{g}\\ \sum_{p^{i}\in W_{obs}^{i}}o_{p^{i}}=\sum_{p^{f}\in W_{obs}^{f}}o_{p^{f}}=\\ \textup{Sum}(O)=\textup{Sum}(O^{\prime}).\end{split}$ (6)
@@ IV-B Collision Cost with Occupancy
We begin by presenting our empirical analysis on how the occupancy change, denoted as $\textup{diff}(O,O^{\prime})$, that results from a ship motion primitive relates to the ship’s kinetic energy loss, collision impulse and approximated work done ($W_{\text{approx}}$ described in Sec. VI-A) during the execution of the motion primitive. For this analysis, we use occupancy maps that match the input/output dimensions of our learned predictor.
@@
$c_{\pi}=\textup{diff}(O,O^{\prime}),$ (8)
@@
$C(\Pi)=\sum_{\pi\in\Pi}\text{diff}(O_{\pi},O^{\prime}_{\pi}),$ (9)
@@ IV-C A* Search with Occupancy Estimation
During each iteration, the current node $n$ is expanded to its neighbor $n^{\prime}$ via a feasible motion primitive $\pi$, denoted by the overloaded $+$ operator at Line 11. A local occupancy $O$ is cropped from the global occupancy $O^{\textup{global}}$ based on the ship pose at node $n$ (Line 13) and is then used for occupancy prediction (Line 14). The edge cost of $\pi$ is then computed as a linear combination of the occupancy-difference collision cost discussed in IV-B scaled by $\alpha$ and the motion primitive length $d(\pi)$ (Line 15). The predicted local occupancy $O^{\prime}$ is stitched back for an updated global occupancy map $O^{\prime\textup{global}}$ (Line 16), which is then cached if the neighbor $n^{\prime}$ is visited for the first time or revisited with a better path (Lines 22). In a future iteration when the node $n^{\prime}$ is being expanded, the global occupancy at $n^{\prime}$ predicted previously is retrieved for further predictions (Line 9). Additionally, by leveraging a GPU in practice, the network inference can be further optimized by performing batched predictions by constructing the inferences from node $n$ to all its neighbors $n^{\prime}$ as a batch.
@@ 18                   Update $n^{\prime}$ parent pointer
IV-D Algorithm Analysis
Here we analyze Alg. 1 runtime and performance.
IV-D1 Runtime
While Alg. 1 has the same structure as $A^{*}$ or Dijkstra’s algorithm, given that the predictor $f$ is a deep neural network, the bottleneck operation is the prediction in Line 14, rather than the priority queue operation in typical cases. Note that in Alg. 1, each node is expanded at most once and is then added to the closed set CLOSED (Line 8). This implies that Alg. 1 runtime scales linearly with the number of edges, or motion primitives (Line 10), explored during the search. Let $|E|$ be the number of edges in the expanded search graph from Alg. 1, it follows that the runtime is $O(|E|P)$, where $P$ is the time for each forward pass of the predictor.
IV-E Performance
We note that by considering obstacle motions, the path finding problem lacks optimal substructures. Namely, in an arbitrary optimal path $n_{s}\leadsto n\leadsto n_{\textup{goal}}$ from start $n_{s}$ to the goal $n_{\textup{goal}}$ that passes through node $n$, the sub-path $n_{s}\leadsto n$ is not necessarily the optimal path from $n_{s}$ to $n$. This could be the cases where, by following the sub-optimal path $n_{s}\leadsto n$ to $n$, the ship pushes the ice floes into a desirable configuration, resulting in the entire path $n_{s}\leadsto n\leadsto n_{\textup{goal}}$ being optimal. Further, optimally solving the navigation among movable obstacle problem is generally considered as NP-hard [24]. Consequently, Alg. 1 lacks theoretical optimality guarantee. Hence, we present the following constant-factor guarantee to upper-bound the performance of Alg. 1.
@@ IV-E Performance
Given a problem instance $I=(n_{s},O_{s},y_{\text{goal}})$, let $\textup{OPT}(I)$ denote the optimal path and let $\textup{Alg}(I)$ denote the path returned by Alg. 1 assuming perfect occupancy predictions. Let $l_{\min}$ be the distance of the shortest action primitive, the cost of $\textup{Alg}(I)$ is upper-bounded as
$u(\textup{Alg}(I))\leq(1+\frac{\alpha}{l_{\min}})\cdot u(\textup{OPT}(I))$ (10)
where $u$ and $\alpha$ are described in Eq. 3.
We present the proof of Theorem. 1 in the next section. We acknowledge that Theorem. 1 may not hold under imperfect predictions, and we leave the incorporation of prediction error uncertainty to future work.
V Algorithm Performance Bound
Recall that given an edge $e=(n_{1},n_{2})$ that connects two nodes $n_{1}$ and $n_{2}$, the cost of the edge $u(e)$ is computed as
$u(e)=d(e)+\alpha\cdot\textup{MSE}(O_{1},O_{2}),$ (11)
where $O_{1}$ and $O_{2}$ are the occupancy maps at nodes $n_{1}$ and $n_{2}$, respectively (Alg. 1, Line 15). This is a linear combination of the edge length $d(e)$ and the occupancy difference from taking the edge. For simplicity we use $u(\cdot)$ to denote both the cost of a path and an edge. By noting that each cell value of an occupancy map $O$ is a value from $[0,1]$, we present the following observation.
Observation 1 (Maximal and Minimal Edge Weights).
For each edge $e$, the edge cost is bounded as $d(e)\leq u(e)\leq d(e)+\alpha$.
Let $I=(n_{s},O_{s},y_{\text{goal}})$ denote an arbitrary problem instance and let $\textup{OPT}(I)=n_{0}\rightarrow n_{1}\rightarrow\ldots\rightarrow n_{k-1}\rightarrow n_{k}$ be the optimal path with $k\geq 1$ edges, where $n_{0}=n_{s}$ and $n_{k}=n_{\textup{goal}}$. We let $e_{i}$ with $i\in\{1,\ldots,k\}$ denote the directed edge on the optimal path from node $n_{i-1}$ to node $n_{i}$.
Recall that in our problem setup the goal is defined to be a line $y_{\textup{goal}}$. So here $n_{k}$ is an arbitrary goal node $n_{\textup{goal}}$ whose y-value is greater than $y_{\textup{goal}}$, and the multiplicity of the goal nodes does not affect our proof. We now prove the following Lemma.
Lemma 1.
For all nodes $n_{i}$ with $i=\{1,\ldots,k\}$ on the optimal path, after running Alg. 1, the cost-so-far for $n_{i}$ found by Alg. 1, denoted as $g(n_{i})$ (Alg. 1 Line 18), is upper-bounded by
$\begin{split}&g(n_{i})\leq d(e_{1})+\ldots+d(e_{i})+\alpha\cdot i.\end{split}$ (12)
Proof.
We prove Lemma 1 by induction on the node $n_{i}$ with $i\in\{1,\ldots,k\}$.
Base Case $n_{i}=n_{1}$: There are two cases:
  1. 1.
    After the execution of Alg. 1, $n_{1}$’s parent pointer points to $n_{0}$. This implies that $g(n_{1})=u(e_{1})\leq d(e_{1})+\alpha$ by Observation 1.
  2. 2.
    After the execution of Alg. 1, $n_{1}$’s parent pointer does not point to $n_{0}$. This implies Alg. 1 finds a cheaper path to $n_{1}$ than following the edge $e_{1}$. It follows that $g(n_{1})\leq u(e_{1})\leq d(e_{1})+\alpha$. Thus, the base case holds.
Inductive Hypothesis: Lemma 1 holds for all nodes $n\in\{n_{1},\ldots,n_{i-1}\}$.
Inductive Step: By Inductive Hypothesis, it follows that after the execution of Alg. 1, we have
$g(n_{i-1})\leq d(e_{1})+\ldots+d(e_{i-1})+\alpha\cdot(i-1).$ (13)
We again proceed with the same case analysis.
  1. 1.
    After the execution of Alg. 1, $n_{i}$’s parent pointer points to $n_{i-1}$. This implies that $g(n_{i})=g(n_{i-1})+u(e_{i})$. By Eq. 13 and Observation 1, we have
    $\begin{split}&g(n_{i})=g(n_{i-1})+u(e_{i})\\ &\leq d(e_{1})+\ldots+d(e_{i-1})+\alpha\cdot(i-1)+d(e_{i})+\alpha\\ &=d(e_{1})+\ldots+d(e_{i})+\alpha\cdot i.\end{split}$ (14)
  2. 2.
    After the execution of Alg. 1, $n_{i}$’s parent pointer does not point to $n_{i-1}$. This implies Alg. 1 finds a cheaper path to $n_{i}$ than following the edge $e_{i}$ from node $n_{i-1}$. Consequently, $g(n_{i})\leq g(n_{i-1})+u(e_{i})$. Applying the same reasoning from Eq. 14, we have
    $\begin{split}&g(n_{i})\leq d(e_{1})+\ldots+d(e_{i})+\alpha\cdot i.\end{split}$ (15)
Thus, by induction, Lemma 1 holds. ∎
With Lemma 1 established, note that $n_{k}=n_{\textup{goal}}$. This implies that after its execution, Alg. 1 is guaranteed to find a path to $n_{\textup{goal}}$ that is cheaper than $d(e_{1})+\ldots+d(e_{k})+\alpha\cdot k$. Let $\textup{Alg}(I)$ denote the path returned by Alg. 1 on instance $I$, it follows that
$\begin{split}&u(\textup{Alg(I)})\leq d(e_{1})+\ldots+d(e_{k})+\alpha\cdot k\\ &=d(\textup{OPT}(I))+\alpha\cdot k.\end{split}$ (16)
Recall that $k$ is the number of edges in the optimal path $\textup{OPT}(I)$, and each edge in Alg. 1 is an action primitive $\pi$. In this work, all distances $d(\cdot)$ are measure in meters. Let $l_{\min}$ denotes the distance of the shortest action primitive in meters. This implies that $k\leq\frac{d(\textup{OPT}(I))}{l_{\min}}$. By expanding $u(\cdot)$ using its definition from Eq. 3, it follows that
$\begin{split}&d(\textup{OPT}(I))+\alpha\cdot k\\ &\leq d(\textup{OPT}(I))+\alpha\cdot k+\alpha\cdot C(\textup{OPT}(I))\\ &\leq d(\textup{OPT}(I))+\frac{\alpha}{l_{\min}}\cdot d(\textup{OPT}(I))+\alpha\cdot C(\textup{OPT}(I))\\ &=(1+\frac{\alpha}{l_{\min}})\cdot d(\textup{OPT}(I))+\alpha\cdot C(\textup{OPT}(I))\\ &\leq(1+\frac{\alpha}{l_{\min}})\cdot(d(\textup{OPT}(I)+\alpha\cdot C(\textup{OPT}(I)))\\ &=(1+\frac{\alpha}{l_{\min}})\cdot u(\textup{OPT}(I)).\end{split}$ (17)
It follows that Theorem 1 holds.
VI Evaluation
@@
VI-A Simulation Setup
We conduct evaluations on the autonomous ship-ice navigation simulator from [3]. The setup is consistent with the experimental platform at the National Research Council Canada ice tank in St. John’s, NL, featuring a 1:45 scale model vessel. Tests are conducted at 20%, 30%, 40%, and 50% concentrations, with 200 randomly generated environments per concentration. Fig. 1 (left) shows a 40% concentration ice field. We evaluate all four planners on the same set of trials, giving a total of $4\,\text{concentrations}\,\times 200\,\text{trials}\,\times 4\,\text{planners}\,=3200\,\text{experiments}$.
@@ VI-B Simulation Results
VI-C Physical Testbed Setup
@@
VI-D Physical Testbed Evaluation
We perform testbed evaluations in 20%, 30%, and 40% concentration environments. For each concentration, 40 trials are performed for each planner, giving a total of $3\,\text{concentrations}\,\times 40\,\text{trials}\,\times 4\,\text{planners}\,=480\,\text{experiments}$. $W_{\text{approx}}$ and travel distance are used as metrics to quantify collision minimization and travel efficiency, respectively. Demonstrations of the testbed setup and experiments can be found in the supplemental video.